### v5

# Decision Machines: An Extension of Decision Trees

Abstract
Here is a compact representation of binary decision trees. We can explicitly draw the dependencies between prediction and binary tests in decision trees and construct a procedure to guide the input instance from the root to its exit leaf. And we provided a connection between decision trees and error-correcting output codes. Then we built a bridge from tree-based models to attention mechanisms.
1 Introduction
The conventional induction of decision tree is greedy to recursively partition the training space. The training space is divided into smaller and smaller subsets to minimize the split criteria until a stopping criterion is satisfied for some training data. As a result, it can be represented graphically as a tree. The leaf nodes take the mode of its ground truth set as the prediction value. It is greedy and data-driven. However, the trained decision tree is always implemented as a collection of ‘if-else-then’ sentences. In another word, it is rule-based, which makes it friendly for us to understand and interpret. In contrast, this implementation is unfriendly for computers to execute. The duality, greedy data-driven training and hard-coded rule-based implementation, does not occur in other supervised machine learning methods.
There are some end-to-end training schemes instead of the greedy training methods for decision trees such as [31, 19, 36, 11, 16]. And the implementation of decision trees can be vectorized such as [25, 6, 23, 26] and RapidScorer[37].
It is a long history to find diverse representation of decision trees since last century. In the section 2 of [9], there is an equivalent model of recursive partitioning in the form of an expansion in a set of basis functions (as products of different step functions) by replacing the geometrical concepts of regions and splitting with the arithmetic notions of adding and multiplying. And it is to transform each tree into a well-known polynomial form in [20]. There are also some schemes to translate the decision trees into neural networks such as [14, 34, 2]. The oblique decision trees are induced from the derivatives of neural models [22].
New representation form would bring or require new optimization and implementation methods. We will show the possibility to reuse the deep learning tool kits to speed up decision trees. Our main contributions include:
  • a provably guaranteed algorithm to represent binary trees in analytic form;
  • a new interpretation of binary decision trees as error correcting output codes;
  • a strong evidence on the possibility to derivate attention mechanisms from decision trees.
2 Logical Decision Machines
In this section, we will show how to represent the binary decision trees in the language of matrix computation. Our main idea is that the true decision path is the maximum likelihood path over all possible paths. If two input instances arrive at different leaf nodes, they must take different paths and there is at least one different results of the Boolean tests. If they arrive at the same leaf node, they must take the same path, where they must share the same results of all tests. We are looking for an encoding scheme to embed the test results and each decision path into a vector space.
In the rest of this section, we will show how to embed the Boolean tests into a binary space and perform an interleaved traversal of decision trees in the means of matrix computation. And we will provide a specific example and some theoretical insights to the nature of decision trees. Finally, we will prove the guarantee of our algorithm and build a connection between decision trees and error-correcting output codes.
2.1 Construction of Logical Decision Machines
We begin with the description of binary decision trees from [25].
Each decision tree $T=(\mathcal{N},\mathcal{L})$ is composed of a set of internal nodes $\mathcal{N}$, and a set of leaves $\mathcal{L}$. Each $n\in\mathcal{N}$ is associated with a Boolean test over a specific feature. Every leaf $l\in\mathcal{L}$ stores the prediction. Considering the test result of a certain internal node $n\in\mathcal{N}$, it is able to infer that some leaves cannot be the exit leaf. Indeed, if $n$ is a false node (i.e., its test condition is false), the leaves in the left subtree of $n$ cannot be the exit leaf. Similarly, if $n$ is a true node, the leaves in the right subtree of $n$ cannot be the exit leaf. Once all the nodes have been processed, the only leaf left is exactly the exit leaf.
For simplicity, we only consider the numerical variables to test, i.e., the input instance $x$ is a real vector in $\mathbb{R}^{n}$. And Boolean conditions are used to test whether the numerical variables are no greater than some given specific thresholds as shown in (1).
In order to represent the true conditions as $-1$s, we substitute 00 with $-1$ in the sign function when $x=0$. And we define the $\operatorname{sgn}(x)$ for $x\in\mathbb{R}$ in our context by
$\operatorname{sgn}(x)=\begin{cases}-1,&\text{if $x\leq 0$},\\ +1,&\text{if $x>0$}.\end{cases}$ (1)
We use the hidden layer to generate the results of all binary tests in decision tree defined by
$h(x)=\operatorname{sgn}(Sx-t),$ (2)
the matrix $S$ is of $\{0,1\}^{(L-1)\times n}$ acting as variable selection to test; the instance $x$ is a real vector in $\mathbb{R}^{n}$; the parameter $t$ is a real vector in $\mathbb{R}^{L-1}$; the integer $L$ is the total number of the leaves.
[…]
Figure 1: A Binary Decision Tree
It is the results rather than the execution order of the Boolean tests that determine the prediction value of the decision trees for a specific input. If we have known all the results of the tests, we want to directly guide the input from the root to the exit leaf.
The template vector of a leaf node in the decision tree is defined by
  1. 1.
    each element corresponds to a distinct internal node;
  2. 2.
    if the leaf is in the right branch of some ancestor node, the element attached to this ancestor is set to $1$;
  3. 3.
    if the leaf is in the left branch of some ancestor node, the element attached to this ancestor is set to $-1$;
  4. 4.
    if an internal node is not a ancestor of the leaf, the element attached to this ancestor is set to 00.
For example, the template vector of the leaf node $v_{3}$ is $(-1,-1,0,+1,+1)$ in the example 1. The template vectors are distinct from each other because each path is unique in decision trees.
If the input takes all the tests, it is supposed to follow the ‘maximum likelihood path’. The key step is to compare the similarities between the result vector $h\in\{1,-1\}^{L-1}$ with the template vectors in $\{1,0,-1\}^{L-1}$.
We modify the cosine similarity into so-called logical similarity defined as
$LS(b,h)=\frac{b\cdot h}{\|b\|_{2}^{2}},$ (3)
where the positive integer $\|b\|_{2}^{2}$ is the squared $\ell_{2}$ norm of template vector $b\in\{1,0,-1\}^{L-1}$. Our similarity(3) is is asymmetric and bounded by $(-1,1]$.
This scheme would enjoy the following properties:
  • Each leaf node must be encoded as a vector of fixed length uniquely.
  • If instance $x$ can reach the leaf node $i$ with template vector $b_{i}$, then the similarity $LS(b_{i},h(x))$ is equal to $1$.
  • If instance $x$ can reach the leaf node $i$ with template vector $b_{i}$, then the similarity $LS(b_{i},h(x))$ is the maximum of all possible paths.
For a binary decision tree, we define its template matrix denoted $B$ in $\{1,0,-1\}^{L\times(L-1)}$ by all the template vectors as its rows. We compute the similarity function (3) on a set of normalized template vectors simultaneously, packed together into a matrix $\bar{B}$ as shown below
$\begin{split}\bar{B}\,h(x))&=(\frac{B_{1}h(x)}{\|B_{1}\|_{2}^{2}},\cdots,\frac{B_{L}h(x)}{\|B_{L}\|_{2}^{2}})^{T}\\ &=(LS(B_{1},h),\cdots,LS(B_{L},h))^{T}\end{split}$ (4)
where the template vector $B_{i}\in\{1,0,-1\}^{L-1}$ is the $i$th row of matrix $B$; the positive integer $\|B_{i}\|_{2}^{2}$ is the squared $\ell_{2}$ norm of the vector $B_{i}$ for $i=1,2,\cdots,L$; the operator $LS(\cdot,\cdot)$ is the logical similarity defined in (3).
The decision trees take the mode of the ground truth within each leaf node as the output prediction. In another word, it is the leaf node that determines the prediction rather than the internal nodes.
We combine (2) and (4) and perform an interleaved traversal of binary trees in the means of matrix computation by
$\begin{split}h&=\operatorname{sgn}(Sx-t)\in\{-1,1\}^{L-1},\\ p&=\bar{B}h\in\mathbb{R}^{L},\\ v[i]&=v[\arg\max{(p)}],\end{split}$ (5)
where the function $\arg\max(p)$ returns the index of maximums in the vector $p$; the index $i$ is equal to $\arg\max(p)$; the output $v[i]$ is the value of the $i$-th leaf node for $i=1,2,\cdots,L$. We call the new formulation of binary decision trees logical decision machines.
Here we do not specify the data type or content form on the value of each leaf. For example, the element $v[i]$ can linear function. Usually it is numerical for regression and categorical/discrete for classification [^1] See more differences of classification trees and regression trees in [24].. And we can embed the class labels into row vectors to make a value matrix $V$, where each column is the class label embedding vector associated with some leaf for classification.
2.2 An Example of the Logical Decision Machine
Suppose the input variable $x$ has 4 numerical features and the trained decision tree is as shown in the figure 1, we will take a specific example to show how logical decision machine(5) works.
The threshold vector is $t=(1,4,3,2,5)^{T}$. The value vector may be categorical or numerical, such as $v=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6})^{T}$. The selection matrix and template matrix of the tree 1 are
$S=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix},\quad B=\begin{pmatrix}-1&-1&0&-1&0\\ -1&-1&0&+1&-1\\ -1&-1&0&+1&+1\\ -1&+1&0&0&0\\ +1&0&-1&0&0\\ +1&0&+1&0&0\\ \end{pmatrix}.$
Suppose the input instance $x$ is $(2,1,2,2)^{T}$, its prediction is $v_{5}$ according to the figure 1. The computation procedure of logical decision tree is as shown below
$Sx-t=\begin{pmatrix}1\\ -3\\ -1\\ -1\\ -3\end{pmatrix},\quad h=\operatorname{sgn}(Sx-t)=\begin{pmatrix}+1\\ -1\\ -1\\ -1\\ -1\end{pmatrix},$
$p=\operatorname{diag}(\|B_{1}\|_{1}^{-1},\cdots,\|B_{L}\|_{1}^{-1})Bh=\begin{pmatrix}\frac{1}{3}\\ 0\\ -\frac{1}{2}\\ 0\\ 1\\ 0\end{pmatrix},$
$i=\arg\max(p)=5,v[i]=v_{5}.$
In summary, we obtain $T(x)=v[\arg\max(\bar{B}\,\operatorname{sgn}(Sx-t))]=v_{5}$.
2.3 Template Matrix
The template matrix $B$ in (5) is related with the structure of decision tree.
The sibling leaves share the same parent and ancestor nodes thus they share the same tests and their unique different result is returned by their parent node. From the computation perspective, the difference of sibling template vectors is parallel to certain one-hot vector.
Each leaf node is attached to a row vector in the template matrix $B$. According to definition of template vector, each internal node is attached to a column vector in the template matrix $B$, where its element is attached to a leaf node. The root node will test all samples so that its corresponding results are not zero in all template vectors. In another word, there is a column vector without any 0 in $B$. For example, the first column of $B$ of the decision tree shown in 1 is associated with the test of the root, which means the test is true for the first four leaves and false for the last two ones. Thus, it implies that the first four leaves belongs to left subtree of the root node and the last two ones belong to the right subtree of the root node. Every test will produce two results, so each column must have at least one $1$ and $-1$. In another word, each test must be involved with at least two decision paths.
Theorem 1
The template matrix $B$ is of full column-rank.
Proof 2.1
It is equivalent to prove that the column vectors of $B$ are linearly independent.
For simplicity, we call the column vectors of $B$ in the right subtree as right vectors and the column vectors of $B$ in the left subtree as left vectors.
The right vectors are independent of the ancestors of the left leaves, so the corresponding elements of left nodes in the right vectors are zero. Thus, those zeroes can lead to the linear independence of the right vectors and the root vector. Similarly, we can prove the linear independence of the left vectors and root vector.
We apply the above procedure to all the subtrees and complete the proof.
It is not difficult to recover the hierarchical structure of certain binary tree if we have known its template matrix. The template matrix is agnostic with respect to the test at the internal branch and the prediction at the leaf node.
Additionally, we can find the subtree is associated with the sub-matrix of the template matrix. For example, we can find the template matrix of the left subtree in 1 as following
$\begin{pmatrix}-1&-1&0\\ -1&+1&-1\\ -1&+1&+1\\ +1&0&0\\ \end{pmatrix}.$
2.4 Provable Guarantee
Now we prove the correctness of the algorithms (5) and build a connection between decision trees and error-correcting output codes.
Theorem 2
The conversion algorithms (5) is correct.
Proof 2.2
We want to prove that $LS(b_{i},h(x))=1$ if and only if the instance $x$ can reach the leaf $i$ where the vector $b_{i}$ is the $i$th template vector of a decision tree for $i\in\{1,\cdots,L\}$ and $h(x)$ is the result vector for an input instance $x$,
If the sample $x$ reach the leaf node $i$, it is easy to verify that $LS(b_{i},h(x))=\max_{i\in\{1,2,\cdots,L\}}LS(b_{i},h(x))=1$.
If $LS(b_{i},h(x))=1$, we can obtain $b_{i},h(x)-b_{i}\right>=0$ by its definition (3) so $h(x)$ is equal to $b_{i}$ for every non-zero element in $b_{i}$, which means the instance $x$ can satisfy the conditions of the $i$th leaf’s ancestor nodes. And the maxima in (4) is unique because $LS(b_{i},h(x))\neq LS(b_{j},h(x))$ when $i\neq j$. Thus, the maxima in (4) is sufficient to identify the exit leaf. And we complete the proof.
And we can obtain the convex-concave upper bound on empirical loss proved in [31] and use the non-greedy methods therein to train the logical decision machines.
We call this template matching interpretation of decision trees because it is to select the prediction value based on similarity between template vectors and the result vector of instances.
The algorithm 5 is also similar to the error correcting output codes (ECOC) such as [7, 18, 1], which is a powerful tool to deal with multi-class categorization problems. The basic idea of ternary ECOC is to associate each class $r$ with a row of a coding matrix $M\in\{-1,0,+1\}^{k\times\ell}$. We briefly introduce ECOC in the algorithms(1).
Algorithm 1 Ternary Error Correcting Output Codes
  1. 1.
    Train the binary classifier $f_{s}$ for each column $s=1,\cdots,\ell$.
  2. 2.
    For training classifier $s$, example labeled $y$ is mapped to $M(y,s)$. Omit examples for which $M(y,s)=0$.
  3. 3.
    Given test example $x$, choose the row $y$ of $M$ that is “closest” to binary predictions $(f_{1}(x),\dots,f_{\ell}(x))$ according to some distance (e.g. modified Hamming).
The bit-vector corresponds to the error-correcting code $h$ and the bit-vector matrix $B$ corresponds to the error-correcting code matrix $M$. All internal nodes are associated with binary test, which are parameterized by $S$ and $t$. And each binary test corresponds to one dichotomizer. In fact, there is a ternary variant of the error-correcting output codes such as [1]. The main difference between 1 and (5) is that the row in the code matrix $M$ corresponds to a specific class while the row in the template matrix $B$ corresponds to a specific leaf. In one hand, we can consider our method as another coding strategy of ternary error-correcting output codes. In another hand, the coding strategy such as [8] will help to find new induction methods of decision tree.
Their main difference is that each class is attached with a single code word in ECOC while each class may be associated with multiple leaves in decision trees.
3 Analytical Decision Machines
To show how the probability concentrates on the exit leaf, we reformulate (5) into three equations:
$\begin{split}h&=\bar{B}\operatorname{sgn}(Sx-t),\\ p&=\arg\max_{d\in\Delta}\left<d,h\right>,\\ T(x)&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (6)
where the notation $\Delta$ is the probability simplex defined as $\Delta=\{d\mid d\geq 0,d^{T}\vec{1}=1\}\subset\mathbb{R}^{L}$; the optima $p$ is in $\{0,1\}^{L}$ where its only non-zero element is in the position of maximum in $h\in\mathbb{R}^{L}$; the integer $L$ is the total number of leaves. The second step in (6) is the key to identify the exit leaf by find the most similar choice. Clearly, the logical decision machine (5) is of hard attention mechanism[35].
Because the only maximum value is equal to $1$ in the similarities defined by (4), we re-express (5) in the following way
$\begin{split}h&=\bar{B}\operatorname{sgn}(Sx-t),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (7)
where the function $\delta(x)$ is equal to $1$ if $x=0$ and 00 otherwise. And $\delta(\cdot)$ is activated to identify the exit leaf. Obviously, both $\bar{B}$ and $S$ are quite sparse matrices, which may help to understand why sparse or quantized neural networks work. However, the shapes of $\bar{B}$ and $S$ may restrict the design space of logical decision machines. Additionally, the operators $\operatorname{sgn}(\cdot)$, $\delta(\cdot)$ and $\arg\max(\cdot)$ are non-differentiable, which isolates logical decision machines from deep neural networks.
The additive tree such as [3, 4] is in the form of
$F(x)=\sum_{i=1}^{N}w_{i}T_{i}(x)$
where $w_{i}\in\mathbb{R}$ and the tree $T_{i}$ is the instances of (7) for $i=1,2,\cdots,N$. It is simple to evaluate (8) in parallel as the same as the random forest [3]. For example, suppose there are two instances $T_{1}(x)$ and ${T}_{2}(x)$ of 7 as following
$h_{1}=\bar{B}_{1}\operatorname{sgn}(S_{1}x-t_{1}),\,T_{1}(x)=\sum_{i=1}^{L_{1}}\delta(1-h_{1}[i])v_{1}[i],$
$h_{2}=\bar{B}_{2}\operatorname{sgn}(S_{2}x-t_{2}),\,T_{2}(x)=\sum_{i=1}^{L_{2}}\delta(1-h_{2}[i])v_{2}[i],$
we can evaluate them in parallel and obtain the sum of trees $w_{1}T_{1}+w_{2}T_{2}$. In fact, the additive tree $w_{1}T_{1}(x)+w_{2}T_{2}(x)$ is yet another instance of (7) as proven below:
$\begin{split}h&=B\operatorname{sgn}(Sx-t),\\ T(x)&=\sum_{i=1}^{L_{1}+L_{2}}\delta(1-h[i])v[i],\end{split}$ (8)
where $B=\begin{pmatrix}\bar{B}_{1}&0\\ 0&\bar{B}_{2}\end{pmatrix}\in\mathbb{R}^{(L_{1}+L_{2})\times(L_{1}+L_{2}-2)},S=\begin{pmatrix}{S}_{1}\\ {S}_{2}\end{pmatrix}\in\{0,1\}^{(L_{1}+L_{2}-2)\times n},$ $t=\begin{pmatrix}t_{1}\\ t_{2}\end{pmatrix}\in\mathbb{R}^{L_{1}+L_{2}-2},v=\begin{pmatrix}w_{1}v_{1}\\ w_{2}v_{2}\end{pmatrix}\in\mathbb{R}^{L_{1}+L_{2}},x\in\mathbb{R}^{n}$. Here we put the multipliers in the value vectors; the function $\delta(\cdot)$ is activated twice for each input example $x$. By this method, we remove the restriction on the shapes of the matrices $\bar{B}$ and $S$ in (5). We are closer to the key-value attention [29, 33].
In the rest of this section, we will build analytical decision machines to approximate or remove some operators in logical decision machines and take advantages of attention mechanisms to expand the design space of decision machines.
3.1 Construction of Analytical Decision Machines
We begin from presenting a simple relaxation of decision machines(7) and, then, we introduce three crucial refinements to the algorithm (5).
We can relax $\operatorname{sgn}(\cdot)$ and $\delta(\cdot)$ in (7) by standard approaches without any modifications of the rest.
The first refinement is to remove the $\operatorname{sgn}$ (1) in (5) or (7). In QuickScorer [25], the exit leaf is identified by the position of the leftmost bit in the product of the false nodes’ bit-vectors, where a bit-vector is assigned to each branch which represents false leaves as $1$ and true leaves as 00. Similarly, we can identify the exit leaf by the position of the rightmost bit in the product of the true nodes’ bit-vectors if we assign a bit-vector to each branch which represents false leaves as 00 and true leaves as $1$. We can identify the exit leaf by the position of the only maximum in the sum of the bit-vectors’ products generated by both algorithms. It is not hard to combine those procedures to remove the $\operatorname{sgn}$ operator in (7) with a well-designed matrix $K$ as described below
$\begin{split}h&=K(Sx-t),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (9)
where other notation is illustrated in(7). The trouble to remove $\operatorname{sgn}$ in (5) is the probability of zero generated by the affine mapping $Sx-t$.
The second refinement turns the algorithm (9) into a soft decision tree, where instances are directed to the exit leaf with some confidence calculated by a scoring function. We apply differentiable $\operatorname{softmax}(\cdot)$ to smoothen distribution over the leaf values in (9):
$T(x)=\sum_{i=1}^{L}\alpha_{i}v_{i},$ (10)
where $(\alpha,\cdots,\alpha_{L})^{T}=\operatorname{softmax}(\frac{K\,(Sx-t)}{\tau})$; the optional $\tau$ is the temperature hyperparameter in $(0,1)$. There are diverse variants of $\operatorname{softmax}$ such as [13, 38, 15, 21]. Here we focus on the positively correlated tests and define our alternative to the $\operatorname{softmax}$ operator in (10) by
$\sum_{i=1}^{L}\frac{\alpha\,sim(K_{i},q)_{+}+1-\alpha}{\alpha\,S+L(1-\alpha)}v_{i},$
where the vector $q=Sx-t$; the real $sim(K_{i},q)_{+}$ is the positive part of the $sim(K_{i},q)$ defined by $\max(0,\left<K_{i},q\right>)$; the total $S$ is $\sum_{i=1}^{L}sim(K_{i},q)_{+}$; the optional $\alpha$ is a hyperparameter in $(0,1]$. It is a special case of [15].
The third refinement is aimed at realization of the soft decision tree[12]. In [12], each internal node directs instances to its children with probabilities given by a sigmoid gating function. It is equivalent to assign a vector to each branch where left leaves are represented as the sigmoid function $\sigma(\cdot)$; right leaves are represented as $1-\sigma(\cdot)$; and other leaves are represented as $1$. We can construct a matrix $M$ to replace $\bar{B}$ in (5) as shown below
$M_{ij}=\begin{cases}\sigma(B_{ij}(w_{j}^{T}x+b_{j})),&\text{if $B_{ij}\neq 0$,}\\ 1,&\text{otherwise},\end{cases}$ (11)
where the function $\sigma(x)=\frac{1}{1+\exp(-x)}$; the integer $B_{ij}\in\{1,0,-1\}$ is the $j$th element in the $i$th row of template matrix $B$ in (4); the weight $w_{j}\in\mathbb{R}^{n}$ and the bias $b_{j}\in\mathbb{R}$ are the parameters associated with $j$th branch. The probability $p[i]$ for each leaf $i$ is equal to $\prod_{j=1}^{L-1}M_{ij}$, which is the product of all the non-zero elements in the $i$th row of $M$. With some arithmetic tricks, the output $y$ of the soft decision tree [12] is evaluated by
$\begin{split}p&=\exp(\ln(M)\vec{1}),\\ y&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (12)
where the matrix $M$ is defined in (11); the vector $\vec{1}$ is a vector of ones with the same row length of $M$; the vector $p$ is the probability distribution of each leaf, i.e.,$\sum_{i=1}^{L}p[i]=1$.
Matrix can contain diverse hierarchical information as shown in (5), (8), (9) and (11).
3.2 Attention-Inspired Decision Machines
Great success of the Transformer architecture proposed in [33] benefits from integral combination of its components. In this section, we will lend support to the design choices of Transformer and borrow some ideas to expand the design space of decision machines.
Now we turn to the scaled dot-product attention in [33] defined by
$Attention(Q,K,V)=\operatorname{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V,$
where a set of queries are packed together into a matrix $Q$ in order to compute the attention function on these queries simultaneously. We can evaluate the attention output $y_{i}$ for each query $q_{i}$ separately as follows
$y_{i}=\operatorname{softmax}(\frac{q_{i}K^{T}}{\sqrt{d_{k}}})V=\sum_{j=1}^{m}\frac{sim(k_{j},q_{i})}{\sum_{n=1}^{i}sim(k_{n},q_{i})}v_{j},$ (13)
where the term $sim(k_{j},q_{i})=\exp(\frac{\left<k_{j},q_{i}\right>}{\sqrt{d_{k}}})$ represents the similarity between the query $q_{i}$ and the key $k_{j}$; the integer $d_{k}$ is the dimension of the key $q_{i}$. Our algorithm (10) is almost identical to the scaled dot-product attention (13), except for the divided temperature of $\tau$. It is a strong evidence that we can derivate the attention mechanism from the decision tree. We relaxed the binary decision tree in (10) from the perspective of model representation, while the scaled dot-product attention is used to let sequence transduction models dispense with recurrence and convolutions in [33].
As shown before, it is the key matrix $B$ that determines the intrinsic hierarchical structure of (8). We can mask the sub-structure of additive trees by matrix projection. For example, We can mask the second tree in (8):
$\begin{split}h_{B}&=BW^{B}\operatorname{sgn}(Sx-t),\\ T_{B}(x)&=\sum_{i=1}^{L_{1}+L_{2}}\delta(1-h_{B}[i])v[i],\end{split}$
where the matrix $W^{B}=\begin{pmatrix}1_{L_{1}\times(L_{1}-1)}&0\\ 0&0_{L_{2}\times(L_{2}-1)}\end{pmatrix}$; the rest are the same as in (8). Here we projected $B$ of (8) to $\begin{pmatrix}\bar{B}_{1}&0\\ 0&0\end{pmatrix}$. This finding may provide some insights into multi-head attention, which originated in [33] to capture the information from different representation subspaces at different positions
$\begin{split}\text{MultiHead}(Q,K,V)&=\text{Concat}(head_{1},\dots,head_{h})W^{O}\\ \text{where $head_{i}$}&=Attention(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})\end{split}$ (14)
where the queries, keys and values are projected into different subspaces by $QW_{i}^{Q}$, $KW_{i}^{K}$ and $VW_{i}^{V}$, respectively.
We can borrow the ideas from sparse attention [30, 33, 27, 5, 10] or linear transformer [32, 17] to expand the design space of decision machines, which are used to reduce the quadratic complexity of scaled dot-product attention in [33]. The first is to use the sparse alternatives to $\operatorname{softmax}$ such as [28, 30, 27, 5, 10]. We prefer these sparse alternatives to $\operatorname{softmax}$ for the sake of more interpretability. The second is to replace the direct computation of similarity between $K$ and $Q$ with a transformation through a non-linear kernel function in linear transformers such as [32, 17]. Those advances in sparse attention and linear transformer permit an efficient acceleration of decision machines (10).
4 Discussion and Conclusion
We represent the binary decision tree in an analytic form (5). We illustrate this representation with a concrete example and explore some intrinsic properties in theory. Then we prove the equivalence of binary decision trees and two-layer networks with sign activation function 2. And we build bridge between the binary decision tree and the ECOC framework. By explicitly formulating the decision tree, we can obtain the upper bound on empirical loss proved in [31].
We remove the $\operatorname{sgn}$ operator and approximate the $\arg\max$ operator in logical decision machines. Based on the proof of 2 we represent the additive trees in analytic form (8). And we find that the analytic decision machines (10) and (11) are closely related with the attention mechanism.
[…]