### v6

# Decision Machines: Enhanced Decision Trees

Abstract
This paper presents Decision Machines (DMs), an innovative evolution of traditional binary decision trees, which leverages matrix computations to significantly enhance both computational efficiency and interpretability. By explicitly mapping the dependencies between predictions and binary tests within a vector space, DMs offer a streamlined approach to navigating decision paths. We integrate decision trees with kernel methods, ensemble methods and attention mechanisms. The integration of these elements not only bolsters the hierarchical structure of decision trees but also aligns with the computational efficiency of matrix computations. Our work bridges the gap between traditional machine learning algorithms and modern deep learning techniques, providing a novel foundation for further research and application in the field of machine learning.
1 Introduction
Decision trees have long been a cornerstone in the field of machine learning, revered for their simplicity and interpretability. However, their traditional induction methods, which are inherently greedy and data-driven, often lead to a representation that is less than optimal for computational efficiency. The training process recursively partitions the training space, dividing it into increasingly smaller subsets until a stopping criterion is met, resulting in a tree structure where leaf nodes predict based on the majority of their ground truth set. This approach, while effective, culminates in an implementation that is rule-based and less amenable to computational execution.
In contrast, other supervised learning methods have moved towards end-to-end training schemes, which offer more flexibility and efficiency. The quest for a more computationally friendly representation of decision trees has led to various innovations, including vectorized implementations and attempts to translate decision trees into neural networks. There are some end-to-end training schemes instead of the greedy training methods for decision trees such as [31, 19, 37, 11, 16]. And the implementation of decision trees can be vectorized such as [25, 6, 23, 26] and RapidScorer[38]. The oblique decision trees are induced from the derivatives of neural models [22]. In the section 2 of [9], there is an equivalent model of recursive partitioning in the form of an expansion in a set of basis functions (as products of different step functions) by replacing the geometrical concepts of regions and splitting with the arithmetic notions of adding and multiplying. And it is to transform each tree into a polynomial form in [20]. There are also some schemes to translate the decision trees into neural networks such as [14, 35, 2]. This paper proposes Decision Machines (DMs), a paradigm that reframes binary decision trees in matrix computation, thereby improving both efficiency and interpretability.
DMs offer a compact representation of binary decision trees, explicitly drawing dependencies between predictions and binary tests. By constructing an arithmetic procedure, we guide input instances from the root to their respective exit leaves, thereby bridging the gap between the intuitive structure of decision trees and the computational efficiency of matrix computations. This formulation not only enhances the interpretability of decision trees but also paves the way for leveraging deep learning toolkits to expedite their execution.
Furthermore, this paper explores the convergence of decision trees with attention mechanisms, a transformative concept in the realm of deep learning. By establishing a connection between tree-based models and attention mechanisms, we open up new avenues for optimizing decision trees and potentially enhancing their predictive power. Our contributions include a provably correct algorithm for representing binary trees analytically, a novel interpretation of binary decision trees as error-correcting output codes, and evidence supporting the derivation of attention mechanisms from decision trees.
2 Logical Decision Machines
We represent binary decision trees in the language of matrix computation, mapping the true decision path as the maximum likelihood path. The logical decision machines embed Boolean tests into a binary space, enabling an interleaved traversal of decision trees through matrix computation.
2.1 Construction
We begin with the description of binary decision trees from [25].
Each decision tree $T=(\mathcal{N},\mathcal{L})$ is composed of a set of internal nodes $\mathcal{N}$, and a set of leaves $\mathcal{L}$. Each $n\in\mathcal{N}$ is associated with a Boolean test over a specific feature. Every leaf $l\in\mathcal{L}$ stores the prediction. Considering the test result of a certain internal node $n\in\mathcal{N}$, it is able to infer that some leaves cannot be the exit leaf. Indeed, if $n$ is a false node (i.e., its test condition is false), the leaves in the left subtree of $n$ cannot be the exit leaf. Similarly, if $n$ is a true node, the leaves in the right subtree of $n$ cannot be the exit leaf. Once all the nodes have been processed, the only leaf left is exactly the exit leaf.
For simplicity, we only consider the numerical variables to test, i.e., the input instance $x$ is a real vector in $\mathbb{R}^{n}$. And Boolean conditions are used to test whether the numerical variables are no greater than some given specific thresholds as shown in (1).
In order to represent the true conditions as $-1$s, we substitute 00 with $-1$ in the sign function when $x=0$. And we define the $\operatorname{sgn}(x)$ for $x\in\mathbb{R}$ in our context by
$\operatorname{sgn}(x)=\begin{cases}-1,&\text{if $x\leq 0$},\\ +1,&\text{if $x>0$}.\end{cases}$ (1)
We use the hidden layer to generate the results of all binary tests in decision tree defined by
$h(x)=\operatorname{sgn}(Sx-t),$ (2)
the matrix $S$ is of $\{0,1\}^{(L-1)\times n}$ acting as variable selection to test; the instance $x$ is a real vector in $\mathbb{R}^{n}$; the parameter $t$ is a real vector in $\mathbb{R}^{L-1}$; the integer $L$ is the total number of the leaves.
[…]
Figure 1: A Binary Decision Tree
It is the results rather than the execution order of the Boolean tests that determine the prediction value of the decision trees for a specific input. If we have known all the results of the tests, we want to directly guide the input from the root to the exit leaf.
The template vector of a leaf node in the decision tree is defined by
  1. 1.
    each element corresponds to a distinct internal node;
  2. 2.
    if the leaf is in the right branch of some ancestor node, the element attached to this ancestor is set to $1$;
  3. 3.
    if the leaf is in the left branch of some ancestor node, the element attached to this ancestor is set to $-1$;
  4. 4.
    if an internal node is not a ancestor of the leaf, the element attached to this ancestor is set to 00.
For example, the template vector of the leaf node $v_{3}$ is $(-1,-1,0,+1,+1)$ in the example 1. The template vectors are distinct from each other because each path is unique in decision trees.
If the input takes all the tests, it is supposed to follow the ‘maximum likelihood path’. The key step is to compare the similarities between the result vector $h\in\{1,-1\}^{L-1}$ with the template vectors in $\{1,0,-1\}^{L-1}$.
We modify the cosine similarity into so-called logical similarity defined as
$LS(k,h)=\frac{k\cdot h}{\|k\|_{2}^{2}},$ (3)
where the positive integer $\|k\|_{2}^{2}$ is the squared $\ell_{2}$ norm of template vector $k\in\{1,0,-1\}^{L-1}$. Our similarity(3) is asymmetric and bounded by $(-1,1]$.
This scheme would enjoy the following properties:
  • Each leaf node must be encoded as a vector of fixed length uniquely.
  • If instance $x$ can reach the leaf node $i$ with template vector $k_{i}$, then the similarity $LS(k_{i},h(x))$ is equal to $1$.
  • If instance $x$ can reach the leaf node $i$ with template vector $k_{i}$, then the similarity $LS(k_{i},h(x))$ is the maximum of all possible paths.
For a binary decision tree, we define its template matrix denoted $B$ in $\{1,0,-1\}^{L\times(L-1)}$ by all the template vectors as its rows. We compute the similarity function (3) on a set of normalized template vectors simultaneously, packed together into a matrix $\bar{B}$ as shown below
$\begin{split}\bar{B}\,h(x))&=(\frac{B_{1}h(x)}{\|B_{1}\|_{2}^{2}},\cdots,\frac{B_{L}h(x)}{\|B_{L}\|_{2}^{2}})^{T}\\ &=(LS(B_{1},h),\cdots,LS(B_{L},h))^{T}\end{split}$ (4)
where the template vector $B_{i}\in\{1,0,-1\}^{L-1}$ is the $i$th row of matrix $B$; the positive integer $\|B_{i}\|_{2}^{2}$ is the squared $\ell_{2}$ norm of the vector $B_{i}$ for $i=1,2,\cdots,L$; the operator $LS(\cdot,\cdot)$ is the logical similarity defined in (3).
The decision trees take the mode of the ground truth within each leaf node as the output prediction. In another word, it is the leaf node that determines the prediction rather than the internal nodes.
We combine (2) and (4) and perform an interleaved traversal of binary trees in the means of matrix computation by
$\begin{split}h&=\operatorname{sgn}(Sx-t)\in\{-1,1\}^{L-1},\\ p&=\bar{B}h\in\mathbb{R}^{L},\\ v[i]&=v[\arg\max{(p)}],\end{split}$ (5)
where the function $\arg\max(p)$ returns the index of maximums in the vector $p$; the index $i$ is equal to $\arg\max(p)$; the output $v[i]$ is the value of the $i$-th leaf node for $i=1,2,\cdots,L$. We call the new formulation of binary decision trees logical decision machines.
Here we do not specify the data type or content form on the value of each leaf. For example, the element $v[i]$ can linear function. Usually it is numerical for regression and categorical/discrete for classification [^1] See more differences of classification trees and regression trees in [24].. And we can embed the class labels into row vectors to make a value matrix $V$, where each column is the class label embedding vector associated with some leaf for classification.
Suppose the input variable $x$ has 4 numerical features and the trained decision tree is as shown in the figure 1, we will take a specific example to show how logical decision machine(5) works.
The threshold vector is $t=(1,4,3,2,5)^{T}$. The value vector may be categorical or numerical, such as $v=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6})^{T}$. The selection matrix and template matrix of the tree 1 are
$S=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix},\quad B=\begin{pmatrix}-1&-1&0&-1&0\\ -1&-1&0&+1&-1\\ -1&-1&0&+1&+1\\ -1&+1&0&0&0\\ +1&0&-1&0&0\\ +1&0&+1&0&0\\ \end{pmatrix}.$
Suppose the input instance $x$ is $(2,1,2,2)^{T}$, its prediction is $v_{5}$ according to the figure 1. The computation procedure of logical decision tree is as shown below
$Sx-t=\begin{pmatrix}1\\ -3\\ -1\\ -1\\ -3\end{pmatrix},\quad h=\operatorname{sgn}(Sx-t)=\begin{pmatrix}+1\\ -1\\ -1\\ -1\\ -1\end{pmatrix},\quad p=\bar{B}h=\begin{pmatrix}\frac{1}{3}\\ 0\\ -\frac{1}{2}\\ 0\\ 1\\ 0\end{pmatrix},$
$i=\arg\max(p)=5,v[i]=v_{5}.$
In summary, we obtain $T(x)=v[\arg\max(\bar{B}\,\operatorname{sgn}(Sx-t))]=v_{5}$.
2.2 Provable guarantee
The template matrix $B$ in (5) is related with the structure of decision tree.
The sibling leaves share the same parent and ancestor nodes thus they share the same tests and their unique different result is returned by their parent node. From the computation perspective, the difference of sibling template vectors is parallel to certain one-hot vector.
Each leaf node is attached to a row vector in the template matrix $B$. According to definition of template vector, each internal node is attached to a column vector in the template matrix $B$, where its element is attached to a leaf node. The root node will test all samples so that its corresponding results are not zero in all template vectors. In another word, there is a column vector without any 0 in $B$. For example, the first column of $B$ of the decision tree shown in 1 is associated with the test of the root, which means the test is true for the first four leaves and false for the last two ones. Thus, it implies that the first four leaves belongs to left subtree of the root node and the last two ones belong to the right subtree of the root node. Every test will produce two results, so each column must have at least one $1$ and $-1$. In another word, each test must be involved with at least two decision paths.
Theorem 1
The template matrix $B$ is of full column-rank.
Proof 2.1
It is equivalent to prove that the column vectors of $B$ are linearly independent.
For simplicity, we call the column vectors of $B$ in the right subtree as right vectors and the column vectors of $B$ in the left subtree as left vectors.
The right vectors are independent of the ancestors of the left leaves, so the corresponding elements of left nodes in the right vectors are zero. Thus, those zeroes can lead to the linear independence of the right vectors and the root vector. Similarly, we can prove the linear independence of the left vectors and root vector.
We apply the above procedure to all the subtrees and complete the proof.
It is not difficult to recover the hierarchical structure of certain binary tree if we have known its template matrix. The template matrix is agnostic with respect to the test at the internal branch and the prediction at the leaf node.
Additionally, we can find the subtree is associated with the sub-matrix of the template matrix. For example, we can find the template matrix of the left subtree in 1 as following
$\begin{pmatrix}-1&-1&0\\ -1&+1&-1\\ -1&+1&+1\\ +1&0&0\\ \end{pmatrix}.$
Now we prove the correctness of the algorithms (5) and build a connection between decision trees and error-correcting output codes.
Theorem 2
The conversion algorithms (5) is correct.
Proof 2.2
We want to prove that $LS(b_{i},h(x))=1$ if and only if the instance $x$ can reach the leaf $i$ where the vector $b_{i}$ is the $i$th template vector of a decision tree for $i\in\{1,\cdots,L\}$ and $h(x)$ is the result vector for an input instance $x$,
If the sample $x$ reach the leaf node $i$, it is easy to verify that $LS(b_{i},h(x))=\max_{i\in\{1,2,\cdots,L\}}LS(b_{i},h(x))=1$.
If $LS(b_{i},h(x))=1$, we can obtain $b_{i},h(x)-b_{i}\right>=0$ by the definition (3) so $h(x)$ is equal to $b_{i}$ for every non-zero element in $b_{i}$, which means the instance $x$ can satisfy the conditions of the $i$th leaf’s ancestor nodes. And the maxima in (4) is unique because $LS(b_{i},h(x))\neq LS(b_{j},h(x))$ when $i\neq j$. Thus, the maxima in (4) is sufficient to identify the exit leaf. And we complete the proof.
2.3 Consistent interpretation
We can provide some consistent interpretations of the logical decision machines with other machine learning models.
We call the proof 2 the template matching interpretation of decision trees because it is to select the prediction value based on similarity between template vectors and the result vector of instances. We can obtain the convex-concave upper bound on empirical loss proved in [31] and use the non-greedy methods therein to train the logical decision machines.
With the test vectors, we can approximate the prediction value of the $j$th leaf node as a kernel based on random forests estimate in [32]
$\ell_{j}.val=\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{D}}\frac{y_{i}\delta(\|\vec{t},\vec{t}_{i}\|_{\mathrm{P}})}{N_{j}}$ (6)
where the function $\delta(x)$ is equal to $1$ if $x=0$ and 00 otherwise; $\vec{t}$ and $\vec{t}_{i}$ are the test vectors of new instance $\mathbf{x}$ and the $i$-th input instance respectively; the number $N_{j}$ is the total sum of samples falling in the $j$th leaf node; the notation $\|\vec{t},\vec{s}\|_{\mathrm{P}}$ is a distance defined as $\|\mathrm{P}(\vec{t}-\vec{s})\|_{2}$.
The algorithm 5 is also similar to the error correcting output codes (ECOC) such as [7, 18, 1], which is a powerful tool to deal with multi-class categorization problems. The basic idea of ternary ECOC is to associate each class $r$ with a row of a coding matrix $M\in\{-1,0,+1\}^{k\times\ell}$. The encoding stage of ECOC is described in the algorithms(1).
Algorithm 1 Ternary Error Correcting Output Codes
1:Input:
  • test example $\mathbf{x}$.
  • the coding matrix $M\in\{-1,0,+1\}^{k\times\ell}$ for class $1,2,\cdots,k$.
  • the pre-trained binary classifiers $f=\{f_{s}\mid s=1,2,\cdots,\ell\}$ for each column of $M$.
2:Output: the predicted label
3:procedure Encoder($\mathbf{x},f$)
4:     Compute the binary predictions of test example $\mathbf{x}$: $p=(f_{1}(x),\dots,f_{\ell}(x))$
5:     for  row $r\in M$ do $\triangleright$ iterate over the rows
6:         $d_{r}\leftarrow\mathrm{dis}(r,p)$
7:     $j\leftarrow$ index of the minimum in $\{d_{1},d_{2},\cdots,d_{\ell}\}$
8:     return $j$
The bit-vector corresponds to the error-correcting code $h$ and the bit-vector matrix $B$ corresponds to the error-correcting code matrix $M$. All internal nodes are associated with binary test, which are parameterized by $S$ and $t$. And each binary test corresponds to one dichotomizer. In fact, there is a ternary variant of the error-correcting output codes such as [1]. The main difference between 1 and (5) is that the row in the code matrix $M$ corresponds to a specific class while the row in the template matrix $B$ corresponds to a specific leaf. In one hand, we can consider our method as another coding strategy of ternary error-correcting output codes. In another hand, the coding strategy such as [8] will help to find new induction methods of decision tree. Their main difference is that each class is attached with a single code word in ECOC while each class may be associated with multiple leaf nodes in decision trees.
3 Analytical Decision Machines
To elucidate the probability concentration on the exit leaf, we have reformulated the process (5) into a more analytical framework, delineating three key equations that capture the essence of our approach.
$\begin{split}h&=\mathrm{\bar{B}}\operatorname{sgn}(Sx-t),\\ p&=\arg\max_{d\in\Delta}\left<d,h\right>,\\ T(x)&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (7)
where the notation $\Delta$ is the probability simplex defined as $\Delta=\{d\mid d\geq 0,d^{T}\vec{1}=1\}\subset\mathbb{R}^{L}$; the optima $p$ is in $\{0,1\}^{L}$ where its only non-zero element is in the position of maximum in $h\in\mathbb{R}^{L}$; the integer $L$ is the total number of leaves. Initially, the hidden decision $h$ for an input instance $x$ is calculated through the feature selection matrix $S$ and the threshold vector $t$. Subsequently, the most probable exit leaf is determined by maximizing the dot product within the probability simplex $\Delta$. Finally, the ultimate prediction $T(x)$ is derived as a weighted sum of the values associated with each leaf, reflecting their respective probabilities. Clearly, the logical decision machine (5) is of hard attention mechanism[36].
Because the only maximum value is equal to $1$ in the similarities defined by (4), we re-express (5) in the following way
$\begin{split}h&=\mathrm{\bar{B}}\operatorname{sgn}(\mathrm{S}x-t),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (8)
where the function $\delta(x)$ is equal to $1$ if $x=0$ and 00 otherwise. And $\delta(\cdot)$ is activated to identify the exit leaf. Obviously, both $\mathrm{\bar{B}}$ and $\mathrm{S}$ are quite sparse matrices, which may help to understand why sparse or quantized neural networks work. However, the shapes of $\mathrm{\bar{B}}$ and $S$ may restrict the design space of logical decision machines. Additionally, the operators $\operatorname{sgn}(\cdot)$, $\delta(\cdot)$ and $\arg\max(\cdot)$ are non-differentiable, which isolates logical decision machines from deep neural networks.
The additive tree such as [3, 4] is in the form of
$F(x)=\sum_{i=1}^{N}w_{i}T_{i}(x)$
where $w_{i}\in\mathbb{R}$ and the tree $T_{i}$ is the instances of (8) for $i=1,2,\cdots,N$. It is simple to evaluate (9) in parallel as the same as the random forest [3]. For example, suppose there are two instances $T_{1}(x)$ and ${T}_{2}(x)$ of 8 as following
$h_{1}=\bar{B}_{1}\operatorname{sgn}(S_{1}x-t_{1}),\,T_{1}(x)=\sum_{i=1}^{L_{1}}\delta(1-h_{1}[i])v_{1}[i],$
$h_{2}=\bar{B}_{2}\operatorname{sgn}(S_{2}x-t_{2}),\,T_{2}(x)=\sum_{i=1}^{L_{2}}\delta(1-h_{2}[i])v_{2}[i],$
we can evaluate them in parallel and obtain the sum of trees $w_{1}T_{1}+w_{2}T_{2}$. In fact, the additive tree $w_{1}T_{1}(x)+w_{2}T_{2}(x)$ is yet another instance of (8) as proven below:
$\begin{split}h&=B\operatorname{sgn}(Sx-t),\\ T(x)&=\sum_{i=1}^{L_{1}+L_{2}}\delta(1-h[i])v[i],\end{split}$ (9)
where $B=\begin{pmatrix}\bar{B}_{1}&0\\ 0&\bar{B}_{2}\end{pmatrix}\in\mathbb{R}^{(L_{1}+L_{2})\times(L_{1}+L_{2}-2)},S=\begin{pmatrix}{S}_{1}\\ {S}_{2}\end{pmatrix}\in\{0,1\}^{(L_{1}+L_{2}-2)\times n},$ $t=\begin{pmatrix}t_{1}\\ t_{2}\end{pmatrix}\in\mathbb{R}^{L_{1}+L_{2}-2},v=\begin{pmatrix}w_{1}v_{1}\\ w_{2}v_{2}\end{pmatrix}\in\mathbb{R}^{L_{1}+L_{2}},x\in\mathbb{R}^{n}$. Here we put the multipliers in the value vectors; the function $\delta(\cdot)$ is activated twice for each input example $x$. By this method, we remove the restriction on the shapes of the matrices $\bar{B}$ and $S$ in (5). We are closer to the key-value attention [29, 34].
In the rest of this section, we will build analytical decision machines to approximate or remove some operators in logical decision machines and take advantages of attention mechanisms to expand the design space of decision machines.
3.1 Construction
We begin from presenting a simple relaxation of decision machines(8) and, then, we introduce three crucial refinements to the algorithm (5).
We can relax $\operatorname{sgn}(\cdot)$ and $\delta(\cdot)$ in (8) by standard approaches without any modifications of the rest.
The first refinement is to remove the $\operatorname{sgn}$ (1) in (5) or (8). The exit leaf is identified by the position of the leftmost bit set to $1$ in the product of the false nodes’ bit-vectors , where are assigned to the internal nodes to represents false leaves as $1$ and true leaves as 00 in QuickScorer [25]. We can combine those procedures and remove the $\operatorname{sgn}$ operator in (8) with well-designed matrices $F$ as described below
$\begin{split}h&=F\operatorname{ReLU}(Wx-b),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (10)
where the operator $\operatorname{ReLU}(Wx-b)$ is defined as $\max(0,Wx-b)$; other notation is illustrated in(8). The trouble to remove $\operatorname{sgn}$ in (5) is the probability of zero generated by the affine mapping $Wx-t$.
The second refinement turns the algorithm (10) into a soft decision tree, where instances are directed to the exit leaf with some confidence calculated by a scoring function. We apply differentiable $\operatorname{softmax}(\cdot)$ to smoothen distribution over the leaf values in (10):
$T(x)=\sum_{i=1}^{L}\alpha_{i}v_{i},$ (11)
where $(\alpha,\cdots,\alpha_{L})^{T}=\operatorname{softmax}(\frac{K\,(Sx-t)}{\tau})$; the optional $\tau$ is the temperature hyperparameter in $(0,1)$. There are diverse variants of $\operatorname{softmax}$ such as [13, 39, 15, 21]. Here we focus on the positively correlated tests and define our alternative to the $\operatorname{softmax}$ operator in (11) by
$\sum_{i=1}^{L}\frac{\alpha\,sim(K_{i},q)_{+}+1-\alpha}{\alpha\,S+L(1-\alpha)}v_{i},$
where the vector $q=Sx-t$; the real $sim(K_{i},q)_{+}$ is the positive part of the $sim(K_{i},q)$ defined by $\max(0,\left<K_{i},q\right>)$; $S$ is $\sum_{i=1}^{L}sim(K_{i},q)_{+}$; the optional $\alpha$ is a hyperparameter in $(0,1]$. It is a special case of [15].
The third refinement is aimed at realization of the soft decision tree[12]. In [12], each internal node directs instances to its children with probabilities given by a sigmoid gating function. It is equivalent to assign a vector to each branch where left leaves are represented as the sigmoid function $\sigma(\cdot)$; right leaves are represented as $1-\sigma(\cdot)$; and other leaves are represented as $1$. We can derive a matrix $M$ from the template matrix $B$ to replace $\bar{B}$ in (5) as shown below
$M_{ij}=\begin{cases}\sigma(B_{ij}(w_{j}^{T}x+b_{j})),&\text{if $B_{ij}\neq 0$,}\\ 1,&\text{otherwise},\end{cases}$ (12)
where the function $\sigma(x)=\frac{1}{1+\exp(-x)}$; the integer $B_{ij}\in\{1,0,-1\}$ is the $j$th element in the $i$th row of template matrix $B$ in (4); the weight $w_{j}\in\mathbb{R}^{n}$ and the bias $b_{j}\in\mathbb{R}$ are the parameters associated with $j$th branch. The probability $p[i]$ for each leaf $i$ is equal to $\prod_{j=1}^{L-1}M_{ij}$, which is the product of all the non-zero elements in the $i$th row of $M$. With some arithmetic tricks, the output $y$ of the soft decision tree [12] is evaluated by
$\begin{split}p&=\exp(\ln(M)\vec{1}),\\ y&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (13)
where the matrix $M$ is defined in (12); the vector $\vec{1}$ is a vector of ones with the same row length of $M$; the vector $p$ is the probability distribution of each leaf, i.e.,$\sum_{i=1}^{L}p[i]=1$.
Matrices in (5), (9), (10) and (12) contain diverse hierarchical information.
3.2 Attention-Inspired models
Great success of the Transformer architecture proposed in [34] benefits from integral combination of its components. In this section, we will lend some support to the design choices of Transformer and borrow some ideas from them to expand the design space of decision machines.
Now we turn to the scaled dot-product attention in [34] defined by
$Attention(Q,K,V)=\operatorname{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V,$
where a set of queries are packed together into a matrix $Q$ in order to compute the attention function on these queries simultaneously. We can evaluate the attention output $y_{i}$ for each query $q_{i}$ separately as follows
$y_{i}=\operatorname{softmax}(\frac{q_{i}K^{T}}{\sqrt{d_{k}}})V=\sum_{j=1}^{m}\frac{sim(k_{j},q_{i})}{\sum_{n=1}^{i}sim(k_{n},q_{i})}v_{j},$ (14)
where the term $sim(k_{j},q_{i})=\exp(\frac{\left<k_{j},q_{i}\right>}{\sqrt{d_{k}}})$ represents the similarity between the query $q_{i}$ and the key $k_{j}$; the integer $d_{k}$ is the dimension of the key $q_{i}$. Our algorithm (11) is almost identical to the scaled dot-product attention (14), except for the divided temperature of $\tau$. It is a strong evidence that we can derivate the attention mechanism from the decision tree. We relaxed the binary decision tree in (11) from the perspective of model representation, while the scaled dot-product attention is used to let sequence transduction models dispense with recurrence and convolutions in [34].
As shown before, it is the key matrix $B$ that determines the intrinsic hierarchical structure of (9). We can mask the sub-structure of additive trees by matrix projection. For example, We can mask the second tree in (9):
$\begin{split}h_{B}&=BW^{B}\operatorname{sgn}(Sx-t),\\ T_{B}(x)&=\sum_{i=1}^{L_{1}+L_{2}}\delta(1-h_{B}[i])v[i],\end{split}$
where the matrix $W^{B}=\begin{pmatrix}1_{L_{1}\times(L_{1}-1)}&0\\ 0&0_{L_{2}\times(L_{2}-1)}\end{pmatrix}$; the rest are the same as in (9). Here we projected $B$ of (9) to $\begin{pmatrix}\bar{B}_{1}&0\\ 0&0\end{pmatrix}$. This finding may provide some insights into multi-head attention, which originated in [34] to capture the information from different representation subspaces at different positions
$\begin{split}\text{MultiHead}(Q,K,V)&=\text{Concat}(head_{1},\dots,head_{h})W^{O}\\ \text{where $head_{i}$}&=Attention(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})\end{split}$ (15)
where the queries, keys and values are projected into different subspaces by $QW_{i}^{Q}$, $KW_{i}^{K}$ and $VW_{i}^{V}$, respectively.
We can borrow the ideas from sparse attention [30, 34, 27, 5, 10] or linear transformer [33, 17] to expand the design space of decision machines, which are used to reduce the quadratic complexity of scaled dot-product attention in [34]. The first is to use the sparse alternatives to $\operatorname{softmax}$ such as [28, 30, 27, 5, 10]. We prefer these sparse alternatives to $\operatorname{softmax}$ for the sake of more interpretability. The second is to replace the direct computation of similarity between $K$ and $Q$ with a transformation through a non-linear kernel function in linear transformers such as [33, 17]. Those advances in sparse attention and linear transformer permit an efficient acceleration of decision machines (11).
4 Discussion and Conclusion
In this paper, we have introduced Decision Machines (DMs), a novel extension of binary decision trees, which leverage matrix computation to enhance their computational efficiency and interpretability. By embedding the decision paths and test results into a vector space, DMs offer a more streamlined approach to model prediction. Our exploration into the integration of DMs with attention mechanisms has opened new avenues for leveraging deep learning toolkits to accelerate decision tree operations.